OpenStudy (anonymous):

PLEASE HELP!!! If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 5x - y = -6 10x + 4y = -6

OpenStudy (anonymous):

a) 10x + 4(-5x - 6) = -6 b) 10x + 4(5x + 6) = -6 c) 10(-5x - 6) + 4y = -6 d) 10(5x + 6) + 4y = -6